# Is the Symbiotic Recurrent Nova T CrB Late? Recent Photometric Evolution and Comparison with Past Pre-Outburst Behaviour

### v1

@@ 1 Introduction
T Coronae Borealis (T CrB) is one of eleven known RNe in the Milky Way (Darnley, 2021; Schaefer, 2022; Schaefer et al., 2022; Shara et al., 2024) and one of only four hosting a red giant donor (T CrB, RS Oph, V745 Sco, and V3890 Sgr), making it a member of the symbiotic subclass of RNe. Among RNe, it is the brightest in quiescence ($V\sim 9.8$ mag; Schaefer, 2023b). The binary system consists of a massive white dwarf that accretes material from an M4III-type red giant companion, in a circular orbit with a period of 227.6 days (Planquart et al., 2025; Hinkle et al., 2025). The mass of the white dwarf has been estimated to be close to the Chandrasekhar limit ($M_{\mathrm{WD}}\approx 1.37\,M_{\odot}$; Hinkle et al., 2025), while the red giant, with a mass of $\sim 0.69\,M_{\odot}$ and a radius of 65 $R_{\odot}$ (Hinkle et al., 2025) fills its Roche lobe, as evidenced by prominent ellipsoidal variability (with an amplitude of $\sim$0.3 mag in $V$). The system lies at a distance of approximately 890 pc (Bailer-Jones et al., 2021; Gaia Collaboration et al., 2023).
@@
As a result, a global observing effort is currently underway, involving both professional and amateur astronomers, with the aim of capturing the next eruption in unprecedented detail across all accessible wavelengths and with virtually every available instrument. The anticipated brightness and accessibility of the system will also make it a unique opportunity for public engagement in transient astronomy.
@@ Acknowledgements
We acknowledge with thanks the variable star observations from the AAVSO International Database contributed by observers worldwide and used in this research.

### v2

@@ 1 Introduction
T Coronae Borealis (T CrB) is one of eleven known RNe in the Milky Way (Darnley, 2021; Schaefer, 2022; Schaefer et al., 2022; Shara et al., 2024a) and one of only four hosting a red giant donor (T CrB, RS Oph, V745 Sco, and V3890 Sgr), making it a member of the symbiotic subclass of RNe. Among RNe, it is the brightest in quiescence ($V\sim 9.8$ mag; Schaefer, 2023b). The binary system consists of a massive white dwarf that accretes material from an M4III-type red giant companion, in a circular orbit with a period of 227.6 days (Planquart et al., 2025; Hinkle et al., 2025). The mass of the white dwarf has been estimated to be close to the Chandrasekhar limit ($M_{\mathrm{WD}}\approx 1.37\,M_{\odot}$; Hinkle et al., 2025), while the red giant, with a mass of $\sim 0.69\,M_{\odot}$ and a radius of 65 $R_{\odot}$ (Hinkle et al., 2025) fills its Roche lobe, as evidenced by prominent ellipsoidal variability (with an amplitude of $\sim$0.3 mag in $V$). The system lies at a distance of approximately 890 pc (Bailer-Jones et al., 2021; Gaia Collaboration et al., 2023).
@@
As a result, a global observing effort is currently underway, involving both professional and amateur astronomers, with the aim of capturing the next eruption in unprecedented detail across all accessible wavelengths and with virtually every available instrument. This will yield data with a level of precision, cadence, and wavelength coverage that has never been achieved for previous eruptions of T CrB, nor for any other object of this class, opening a unique window for new discoveries and for explaining the peculiarities of T CrB. Among other things, T CrB is predicted to become the brightest classical or recurrent nova ever observed in X-rays, with two expected X-ray episodes (Starrfield et al., 2025), including an early X-ray flash, so far seen only in the classical nova YZ Ret (König et al., 2022). Due to strong shocks, it is also expected to be a prominent source of very high-energy $\gamma$-rays, similar to RS Oph (H. E. S. S. Collaboration et al., 2022; Acciari et al., 2022).
Infrared spectroscopy, including integral field unit spectrographs, can be used to study nucleosynthesis and the structure of the ejecta (see Starrfield et al., 2025), while infrared interferometry will allow spatially resolved studies of the ejecta as it expands (e.g., Otulakowska-Hypka, 2025). High-resolution optical and infrared imaging may also enable the detection of light echoes from the recently discovered nova super-remnant surrounding T CrB (Shara et al., 2024b). While fluorescent light echoes are unlikely to be detectable according to the authors, dust-scattered continuum echoes might be observed, probing the distribution and properties of the circumstellar dust.
It will also be interesting to investigate the impact of the eruption on the orbital period of the system. Schaefer (2023b) reported a significant period change after the 1946 eruption based on photometric data. Now, for the first time, we will have the opportunity to compare precise spectroscopic orbital solutions before and after an eruption. Finally, the anticipated brightness and accessibility of the system will also make it a unique opportunity for public engagement in transient astronomy.
@@ Acknowledgements
We thank the referee, Michael Shara, for a careful reading of the manuscript and for the constructive suggestions that helped improve it. We acknowledge with thanks the variable star observations from the AAVSO International Database contributed by observers worldwide and used in this research.